OpenStudy (anonymous):

PLEASE HELP ME Write the equation of direct variation that includes the given point. (6, 5) (4, -8)

OpenStudy (anteater):

Direct variation functions look like this: y = kx So, they give you an x and a y (since your point is an x,y pair). And you just need to plug them in and figure out what k is.

OpenStudy (e.mccormick):

Also, all direct variation functions go throught he origin, so you have two points.

OpenStudy (anteater):

So, if you have the point (6,5). They y = kx --> 5 = k(6) . Therefore 5/6 = k

OpenStudy (anteater):

And so your direct variation function is: y = (5/6)x

OpenStudy (anteater):

So then, I think they want you to find another direct variation function that contains the point (4,-8)

OpenStudy (anteater):

So what would that direct variation function look like?

OpenStudy (anteater):

And yes, all direct variation functions do go through the origin. :)

OpenStudy (anonymous):

y=-2x?

OpenStudy (e.mccormick):

Yes.

OpenStudy (anteater):

Yep :)